Determine the chemical formulas and molecular weights for the compounds
Answer #1

Ans :-

a. Too CHO 18 26 SoHo Hat is F18 H2O that is Molecular weight = 18x12 + 26X) + 1x TESSO Molecular weight = 13X12 +22x1+3x 16(15H2ool Molecular weight = 15x12 + 20x1+1x 16 =121647 ZO Ice Hi Oy that is C12H14 Oy Molecular kleight = 12x12 +14X144X16 =